# In-and-Out: Algorithmic Diffusion for Sampling Convex Bodies

### v3

@@ Appendix A $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
Direct analysis.
The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of the $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [Vem05, §5].
@@
Indirect analysis through $\mathsf{Speedy\ walk}$.
[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 2), which is slightly biased from $\pi$. Then, the $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, the $\mathsf{Ball\ walk}$ often refers to $\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of the $\mathsf{Ball\ walk}$.
@@ Appendix B Functional inequalities
Cheeger and PI constants.
The Cheeger isoperimetric constant $C_{\mathsf{Ch}}(\mu)$ measures how large surface area a measurable subset with larger volume has, defined by
@@
Log-Cheeger and LSI constants.
Just as the Cheeger and PI constants are related above, there are known connections between LSI and log-Cheeger constants. The log-Cheeger constant $C_{\mathsf{logCh}}(\mu)$ of a distribution $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is defined as
@@ Appendix C Semigroup calculus for contraction under the heat flow
Forward heat flow.
We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
@@
Backward heat flow.
Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.

### v4

@@ Appendix A $\mathsf{Ball\ walk}$ and $\mathsf{Speedy\ walk}$
Direct analysis.  The following $\mathsf{TV}$-guarantee is obtained by lower bounding the $s$-conductance of the $\mathsf{Ball\ walk}$, which requires a one-step coupling argument and the Cheeger inequality for $\pi$. We refer interested readers to [Vem05, §5].
@@
Indirect analysis through $\mathsf{Speedy\ walk}$.[KLS97] introduced the $\mathsf{Speedy\ walk}$, which could be viewed as a version of the $\mathsf{Ball\ walk}$ and converges to a speedy distribution (see Proposition 3), which is slightly biased from $\pi$. Then, the $\mathsf{Speedy\ walk}$ is used together with another algorithmic component (rejection sampling) [KLS97, Algorithm 4.15] that converts the speedy distribution to the uniform distribution. In the literature, the $\mathsf{Ball\ walk}$ often refers to $\mathsf{Speedy\ walk}$ combined with the conversion step’, rather than a direct implementation of Algorithm 2. Strictly speaking, a mixing guarantee of this combined algorithm should not be referred to as a provable guarantee of the $\mathsf{Ball\ walk}$.
@@ Appendix B Functional inequalities
Cheeger and PI constants.  The Cheeger isoperimetric constant $C_{\mathsf{Ch}}(\mu)$ measures how large surface area a measurable subset with larger volume has, defined by
@@
Log-Cheeger and LSI constants.  Just as the Cheeger and PI constants are related above, there are known connections between LSI and log-Cheeger constants. The log-Cheeger constant $C_{\mathsf{logCh}}(\mu)$ of a distribution $\mu\in\mathcal{P}(\mathbb{R}^{d})$ is defined as
@@ Appendix C Semigroup calculus for contraction under the heat flow
Forward heat flow.  We begin by introducing the “heat flow” equation (or also known as the Fokker–Planck equation), which describes the evolution of the law of $Z_{t}$ under ($\mathsf{FH}$),
@@
Backward heat flow.  Although there are many ways to define a “reversal” of $P_{h}$, we will use the notion of adjoint introduced by [KP23], which is the most immediately useful.
@@
Appendix D Bias from failure
By computing the one-step distribution, we can note that $\mathsf{In\text{-}and\text{-}Out}$ conditioned on success has a bias. To see this, consider $\mathsf{In\text{-}and\text{-}Out}$ with threshold $N$ and the initial measure $X_{1}\sim\mu$. Then, for a measurable set $A$ and the success event $S$,
$\mathbb{P}(X_{2}\in A\mid S)$ $=\frac{\mathbb{P}(X_{2}\in A,S)}{\mathbb{P}(S)}=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,X_{1}=x,S)\,\mathrm{d}x$
$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid X_{1}=x)\,\mu(\mathrm{d}x)$
$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S,Y_{2}=y\mid X_{1}=x)\,\mu(\mathrm{d}x)\,\mathrm{d}y$
$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid X_{1}=x,Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(X_{2}\in A,S\mid Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\mathbb{P}(S)}\int\mathbb{P}(S\mid Y_{2}=y)\mathbb{P}(X_{2}\in A\mid S,Y_{2}=y)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\mathbb{P}(S)}\int\bigl(1-\bigl(1-\ell(y)\bigr)^{N}\bigr)\,\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y$
$=\frac{\int\bigl(1-\bigl(1-\ell(y)\bigr)^{N}\bigr)\,\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y}{1-\mathbb{E}_{\mu*\gamma}[(1-\ell)^{N}]}\,.$
On the other hand, $\mathsf{In\text{-}and\text{-}Out}$ without cap (i.e., $N=\infty$) satisfies
$\mathbb{P}(X_{2}\in A)=\int\pi^{X|Y=y}(A)\,\gamma(y-x)\,\mu(x)\,\mathrm{d}x\mathrm{d}y\,.$
Thus, trajectories are weighted differently in the two distributions.